OpenStudy (anonymous):

a forest ranger spots a fire on a 28 foot tower,the angle of depression from the tower to the fire is 11 degrees to the nearest foot how far is the fire from the base of the tower show the steps you use to find the solution running short on time to turn this answer in plz help

OpenStudy (hlambach):

Now, since the angle of depression is 11°, we can know the angle inside the triangle is 79°, since 11+79=90 (Forms right angle). Now, since a right angle is 90°, and since all angles in a triangle sums up to 180° (Triangle Angle Sum Theorem), the angle of elevation is 11°. Now we can use the Law of Sines to find the answer ("x" is the side we are looking for). x/sin(90) = 28/ sin(11) x = (28/sin(11))sin(90) And, using a calculator: x= 147 (rounded to the nearest foot) So the distance between the tower and the fire is 147 feet. Hope this helps!  BTW: Sorry if the drawing is a little confusing :)